Consider the following equations: Qd = 31 - 3P and Qs = 6 + 2P, where Qd = quantity demanded and Qs = quantity supplied, P = price
Demand Schedule
Price Quantity demanded
$250 1500
$200 2100
$150 2700
1
Expert's answer
2019-05-07T10:10:55-0400

Solution: Demand equation ,"Q_d = 31 -3P"  ……………………..(1)

Supply equation, "Q_s = 6 + 2P"   ……………………..(2)

So, in equilibrium, "Q_d = Q_s"

                            or, "31 - 3P = 6 + 2P"

                            or, "5P = 25"

                            or, "P = 5"

 

So, "Q_d = Q_s = 31 \u2013 15 = 16" [As, P = 5]


So, "P = 5" and "Q = 16"

 

 

Answer: So, price is $5 and quantity demanded is 16 unit.